# Graph Neural Network Training Systems: A Performance Comparison of Full-Graph and Mini-Batch. [Experiment, Analysis & Benchmark]

### v1

@@ Abstract.
In this paper, we provide a comprehensive empirical comparison of representative full-graph and mini-batch GNN training systems. We find that the mini-batch training systems consistently converge faster than the full-graph training ones across multiple datasets, GNN models, and system configurations, with speedups between 2.4×2.4$\times$2.4 × - 15.2×15.2$\times$15.2 ×. We also find that both training techniques converge to similar accuracy values. Our work highlights the importance of comparing system across different classes, using time-to-accuracy rather than epoch time for performance comparison, and selecting appropriate hyperparameters for the comparison.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2025.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/saurabhbajaj123/GNN_minibatch_vs_fullbatch.
@@ 1. Introduction
Two classes of GNN systems: Full-graph and mini-batch. GNNs can be trained using either a mini-batch or a full-batch (typically called full-graph) approach, much like other machine learning models. In GNN training, however, these two approaches result in different message-passing patterns, which require different system pipelines and pose different system optimization challenges. This resulted in the development of two distinct classes of GNN training systems, each designed to support scalability for either mini-batch or full-graph training.
@@
The main reason for the emergence of these two separate class of systems is that different training methods (full-graph vs. mini-batch) require different data management pipelines to partition data, transfer it across CPUs and GPUs, and parallelize computation when scaling to large graphs. Full-graph GNN training performs message-passing for the entire batch, typically across the entire graph, at each epoch. To scale to large graphs that do not fit in the memory of a single GPU, multi-GPU full-graph training systems use model parallelism: they partition the graph, process different partitions at different GPUs, and perform message passing across partitions (Tripathy et al., 2020; Ma et al., 2019; Jia et al., 2020). For example, the PipeGCN (Wan et al., 2022b) full-graph training system partitions the input graph and keeps each partition in a different GPU (see Figure 1a). The black dotted lines show the communication of vertex features across GPUs during message passing at each GNN layer in the forward and backward pass. The communication happens through either a fast cross-GPU bus such as NVLink, if available, or PCIe.
@@
Mini-batch training breaks up the training set into mini-batches. Each training epoch now consists of multiple iterations, one per mini-batch. Mini-batch training is popular in production systems because it is amenable to scaling through data parallelism (Zheng et al., 2020, 2022b; Zhu et al., 2019). A unique challenge when applying data parallelism to GNNs is that preparing mini-batches requires expensive pre-processing at each iteration. At each iteration, data-parallel mini-batch training systems sample the k-hop neighborhood of the vertices in the mini-batch, load each sample into a different GPU and perform local message passing only within each sample, without passing message across GPUs. For example, Figure  1b shows a training iteration in the popular DGL system. First, each GPU obtains a sample of the k-hop neighborhood of a subset of the vertices in the mini-batch. Sampling is preferably executed using the GPU since it is a computationally expensive step in itself (Jangda et al., 2021). Next, the sample and the input features of its vertices are loaded to the GPUs. Finally, each GPU performs a forward and backward pass on the local sample independently.
@@
These different data management pipelines result in a different set of optimizations. Full-graph training systems must deal with expensive and irregular computation and high communication costs at each epoch. They propose sophisticated scheduling and partitioning algorithms to address these bottlenecks (Md et al., 2021; Peng et al., 2022; Wan et al., 2022b; Thorpe et al., 2021; Wan et al., 2023a). Work can be partitioned and scheduled offline, since the graph is static. Some full-graph training systems also propose techniques to reduce the volume of message data transferred across GPUs and hosts (Wan et al., 2023b; Peng et al., 2022; Wan et al., 2022a). In data-parallel mini-batch training systems, instead, computation at each iteration is much more lightweight and there is no cross-GPU message passing. However, pre-processing is a major bottleneck. Samples are typically large, so sampling them and loading them into GPUs is expensive. Research on scalable mini-batch systems has focused on addressing these bottlenecks by accelerating mini-batch sampling (Yang et al., 2022b; Jangda et al., 2021; Pandey et al., 2020; Wang et al., 2021a; Gong et al., 2023; Cai et al., 2023; Zheng et al., 2022a, b) and reducing the cost of loading the input features by caching them in the memory of the GPUs (Lin et al., 2020; Yang et al., 2022b; Song et al., 2023b; Cai et al., 2023; Tan et al., 2023; Yang et al., 2022a; Liu et al., 2023; Polisetty et al., 2023).
@@
Motivation: What is the state of the art? To understand the state of the art in the field it is necessary to compare systems across different classes. We performed an exhaustive review of the existing literature and found that prior works that propose systems in one class compare them with other systems in the same class, with limited or no performance evaluation of systems in the other class. Furthermore, several papers assert the superiority of the training approach they support in terms of the best accuracy that can be achieved, resulting in conflicting claims. We report a detailed discussion of the evidence in the literature in Section  2. Overall, the existing literature makes it difficult to get a clear picture of the state of the art of the field.
Performance evaluation. In this paper, we address this open issue by performing a thorough performance comparison of representative state-of-the-art GNN training systems across the two classes (Wan et al., 2022b, a; Wang et al., 2019; Zheng et al., 2020; Tan et al., 2023). We chose systems that have a stable, publicly-available, and open-source implementation, are representative of the most important types of systems optimizations in their class, and implement all the standard GNN models (GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), and GAT (Veličković et al., 2018)) in their repository.
@@
Our results show that the mini-batch training systems are consistently faster than the full-graph training counterparts in reaching the same target accuracy across all datasets, GNN model, and hardware deployments we considered. These results are consistent with the evidence in the literature. Several papers that propose full-graph training systems use epoch time as the main metric to compare with mini-batch training systems, showing that the latter have a larger epoch time  (Wang et al., 2022; Wan et al., 2023a; Md et al., 2021; Chen et al., 2023). Our evaluation confirms these results. However, our results also highlight that considering the epoch time alone only gives a partial picture and that time-to-accuracy is a better metric for comparison. Mini-batch training typically requires fewer epochs to converge because it updates the model parameters multiple times per epoch, once at each iteration, whereas full-graph training performs only one update per epoch. Therefore, even if mini-batch training systems perform more work per epoch than full-graph training ones to run multiple iterations, their overall time-to-accuracy is lower. Our empirical results show that this holds consistently across all systems, deployments, models, and datasets we consider. Across all the datasets we considered, the fastest mini-batch training system is between 2.4×2.4$\times$2.4 × to 15.2×15.2$\times$15.2 × (6.3×6.3$\times$6.3 × on average) faster than the fastest full-graph systems in terms of time-to-accuracy.
In terms of scalability, full-graph training and mini-batch training systems observe similar relative speedups when they scale from a single GPU to multiple GPUs. However, the mini-batch systems we consider have a faster time-to-accuracy on a single GPU or host, so they remain faster in absolute terms in all configurations. We report our performance results in Section 4.
Performance vs. accuracy. The time-to-accuracy experiments discussed previously measure the time it takes different systems to reach the same target accuracy, which we select by performing hyperparameter tuning, using the same model and dataset. However, using a system with a longer time-to-accuracy can be justifiable if a model can converge to a higher accuracy. This is an important factor when navigating the performance trade-offs of using different GNN training systems. In fact, some papers proposing full-graph GNN systems assert that full-graph training achieves higher accuracy (Thorpe et al., 2021; Wan et al., 2022b, a, 2023a; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wan et al., 2023b; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019), while others proposing mini-batch training systems claim that mini-batch training achieves higher accuracy (Kaler et al., 2022; Zou et al., 2019; Zheng et al., 2022b; Zhao et al., 2021; Zheng et al., 2020).
@@
Our results show that, with proper hyperparameter tuning, both full-graph and mini-batch training reach a similar accuracy across different datasets and models. Which method achieves the highest accuracy depends on the dataset and GNN model but the accuracy gaps are small. In our evaluation, we matched or exceeded all the accuracy results in the literature on GNN training systems we reviewed. We observe that given a dataset and GNN model, if we use the best hyperparameters for one training method with a system using the other training method, the latter achieves lower accuracy. This dependence on the choice of hyperparameters may explain the contradicting results found in the literature. Our results underline the importance of conducting separate hyperparameter tuning for the two training methods when comparing performance and accuracy across systems using different training methods. We also find that finding good hyperaparameters is easier with mini-batch training because it is less sensitive to variations in the model architecture. This has an impact on the overall training time, since model training often involves performing hyperparameter search. The harder it is to find good hyperparameters, the longer the overall training time. We report our accuracy experiments in Section 5.
@@
Lessons learned. Our results show that the mini-batch training systems we consider consistently achieve a better performance than the full-graph training ones and a similar accuracy. We stress, however, that the goal of this study is not to establish the inherent superiority of one class of systems over the other. Our study is an empirical evaluation and the field is continuously evolving.
From a methodological perspective, our work highlights the importance of comparing GNN training systems across different classes. It also indicates a set of principles on how to perform this comparison: use time-to-accuracy as primary metric for comparison rather than epoch time, and perform separate hyperparameter tuning for mini-batch and full-graph training when selecting a target accuracy for performance comparison or comparing convergence accuracy. Prior works do not always follow these principles, and this may explain the apparent contradiction in the existing results. Our work offers guidance to ML practitioners who use GNN training systems and to systems researchers who design and evaluate them.
An extended version of this paper, which includes additional experimental results that we had to omit due to space constraints, is available at (Bajaj et al., 2024). To ease reproducibility, our study uses publicly available datasets, model implementations, and training systems. We make all our scripts publicly available. [^1] https://github.com/saurabhbajaj123/GNN_minibatch_vs_fullbatch
@@ 2. Background and Motivation
This section provides necessary background on GNN and GNN training systems to follow the rest of the paper. It then discusses the existing evidence in the literature.
@@ 2.1. Graph Neural Networks
Given a graph $G(V,E)$, where each node $v\in V$ is represented with a feature vector $h_{v}^{0}$ (input feature of vertex $v$) and the edges between any two vertices $u$ and $v$ is $e_{uv}\in E$. A GNN at $l^{th}$ layer performs the following computation to get the representation of any vector,
@@
where $h_{v}^{l}$ is the $l^{th}$ layer representation for vertex $v$, $N$ represents the incoming neighbor vertices of vertex $v$, $W^{l}$ is the model weight matrix, $\Psi^{l}$ is any aggregation function and $\Phi^{l}$ is an update function.
This paper considers the three standard GNN models used in previous work: GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), and GCN (Kipf and Welling, 2017). GraphSAGE learns the representation of a target node based on the features of the nodes in its k-hop neighborhood. The information from the neighboring nodes is propagated to the target node by performing message passing and aggregation recursively, one hop at a time. GAT (Graph Attention Networks) extend message passing by drawing inspiration from attention architectures in transformer models. It uses an attention score to weigh the incoming message of a node. GCN learns hidden layers representations and encode the graph structure and the features of the nodes by implementing a first-order approximation of spectral convolutions on graphs.
@@
Proposes System/paper Main Claims Experimental evidence Year Open Source Platform
@@
AdGNN (Song et al., 2023a) FG higher ACC, FG better convergence 2023 Multi-host single-GPU
G3 (Wan et al., 2023a) FG higher ACC, lower ET ACC, ET 2023 Multi-host multi-GPU
AdaQP (Wan et al., 2023b) MB information loss 2023 Mutli-host mult-GPU
@@
DGCL (Cai et al., 2021) MB ACC loss 2021 Mutli-host mult-GPU
@@
Full-graph and mini-batch training systems implement different data management pipelines for scalable multi-GPU training, which resulted in the emergence of two separate categories of systems. For a more detailed discussion on distributed GNN training systems, we refer the reader to (Shao et al., 2024; Lin et al., 2023).
Full-graph training systems The key idea behind full-graph training systems is to partition a large graph into multiple subgraphs such that each can fit into a GPU, and then train them in parallel. This requires communication of vertex features across GPUs. Different types of communication strategies have been used in the literature, including broadcast-based (Tripathy et al., 2020), point-to-point based (Md et al., 2021; Demirci et al., 2022), pipeline-based (Wan et al., 2023a; Mostafa, 2022; Md et al., 2021; Wang et al., 2021b), and communication via shared memory (Jia et al., 2020; Ma et al., 2019).
Communication incurs a substantial performance overhead and requires frequent synchronization. To overcome this bottleneck, two main types of optimizations have been proposed: asynchronous communication and sampling. Asynchronous communication allows the use of stale local information to avoid expensive synchronization costs. They can be categorized as epoch-fixed (Md et al., 2021; Wan et al., 2022b) and epoch-adaptive (Thorpe et al., 2021) staleness models. In GNN training, two synchronization points exist which are specific to GNN training: first, embedding aggregation and second, gradient aggregation. Some protocols (Peng et al., 2022; Md et al., 2021) only consider the embedding aggregation stage to support asynchronous aggregation, while others (Wan et al., 2022b; Thorpe et al., 2021) are designed to use historical information in both the embedding aggregation and gradient aggregation stages. Sampling keeps training synchronous but reduces communication cost by exchanging messages only among a subset of the boundary vertices across partitions (Wan et al., 2022a). Synchronous training is preferred in many cases because its results are predictable and reproducible.
Mini-batch training systems Mini-batch training systems rely on the generation of mini-batches to scale GNNs to large graphs and employ graph sampling to reduce the neighborhood explosion problem. Mini-batch training systems have a different set of performance bottlenecks than full-graph training ones. Existing research has focused on optimizing two steps, mini-batch sampling and input feature loading, which are major performance bottlenecks and often have a cost comparable to training. Multiple vertex-wise, layer-wise and subgraph-wise sampling techniques have been proposed (Chen et al., 2018a; Hamilton et al., 2017; Zeng et al., 2019; Zou et al., 2019). It is common to accelerate these algorithms using GPUs (Jangda et al., 2021; Pandey et al., 2020; Wang et al., 2021a; Gong et al., 2023; Yang et al., 2022b). To avoid massive CPU-GPU communication while loading, some works (Zhu et al., 2019) have proposed caching techniques, like replicating frequently accessed remote vertices at local , store high out-degree vertices (Lin et al., 2020), caching hot vertices (i.e., those accessed more often) (Min et al., 2022). Some works (Liu et al., 2023) propose dynamic caching using a FIFO policy. Some distributed sampling techniques (Jiang and Rumi, 2021; Cai et al., 2023) prioritize local vertices, pushing sampling tasks to remote workers only if results need to be sent back, reducing communication. Some cache both topological and feature data on the GPUs (Sun et al., 2023; Zhang et al., 2023; Cai et al., 2023). To avoid graph sampling becoming a bottleneck, some works have focused on developing efficient GNN dataloaders (Bai et al., 2021; Deutsch et al., 2019; Dong et al., 2021; Min et al., 2021). Some systems reduce the cost of loading by proposing hybrid parallelism approaches for mini-batch training that are alternative to data parallelism: push-pull parallelism (Gandhi and Iyer, 2021) and split parallelism (Polisetty et al., 2023).

### v3

@@ Abstract.
In this paper, we provide a comprehensive empirical comparison of representative full-graph and mini-batch GNN training systems. We find that the mini-batch training systems consistently converge faster than the full-graph training ones across multiple datasets, GNN models, and system configurations. We also find that mini-batch training techniques converge to similar or often higher accuracy values as full-graph training ones, showing that mini-batch sampling is not necessarily detrimental to accuracy. Our work highlights the importance of comparing systems across different classes, using time-to-accuracy rather than epoch time for performance comparison, and selecting appropriate hyperparameters for each training method separately.
@@ 1. Introduction
Two classes of GNN systems: Full-graph and mini-batch. GNNs can be trained using either a mini-batch or a full-batch (typically called full-graph) approach, much like other machine learning models. In standard deep neural network (DNN) training, the dataset consists of individual training examples that can be processed independently and have no structural dependencies. In GNNs, in contrast, the training data is composed of vertices that are interconnected through edges, forming a graph structure where vertices cannot be treated as independent training examples. Full-graph and mini-batch training deal with these dependencies with different data management pipelines to partition data and parallelize computation and communication when scaling to large graphs. This resulted in the development of two distinct classes of GNN training systems, each designed to support either mini-batch or full-graph training.
@@
Full-graph GNN training performs message-passing across the entire graph at each epoch. To scale to large graphs that do not fit in the memory of a single GPU, multi-GPU full-graph training systems use model parallelism: they partition the graph, process different partitions at different GPUs, and exchange hidden vertex features across partitions (Tripathy et al., 2020; Ma et al., 2019; Jia et al., 2020). For example, the PipeGCN (Wan et al., 2022b) full-graph training system partitions the input graph and keeps each partition in a different GPU (see Figure 1(a)). The black dotted lines show the communication of hidden vertex features across GPUs at each GNN layer in the forward and backward pass. The communication happens through either a fast cross-GPU bus such as NVLink, if available, or PCIe.
@@
Mini-batch training breaks up the training set into mini-batches. Each training epoch now consists of multiple iterations, one per mini-batch. Mini-batch training is popular in production systems because it is amenable to scaling through data parallelism (Zheng et al., 2020, 2022b; Zhu et al., 2019). A unique challenge when applying data parallelism to GNNs is that preparing mini-batches requires expensive pre-processing at each iteration. At each iteration, data-parallel mini-batch training systems sample a subset of the k-hop neighbors of the vertices in the mini-batch, load each sample into a different GPU, and perform local message passing only within each sample, without passing message across GPUs. For example, Figure  1(b) shows a training iteration in the popular DGL system. First, each GPU obtains a sample for a subset of the vertices in the mini-batch. Graph sampling is preferably executed using the GPU since it is a computationally expensive step in itself (Jangda et al., 2021). Next, the sample and the input features of its vertices are loaded to the GPUs. Finally, each GPU performs a forward and backward pass on the local sample independently.
@@
These different data management pipelines result in a different set of optimizations. Full-graph training systems must deal with expensive and irregular computation and high communication costs at each epoch. In mini-batch training systems, instead, computation at each iteration is much more lightweight and there is no cross-GPU exchange of hidden features at each layer. However, pre-processing is a major bottleneck. Samples are typically large, so sampling them and loading them into GPUs is expensive.
The common rationale for using full-graph GNN training is to ensure that the training algorithm sees all the required dependent data in the graph, which is assumed to be the key to achieving high accuracy (Thorpe et al., 2021; Wan et al., 2022b, a, 2023b; Tripathy et al., 2020; Cai et al., 2021; Yang et al., 2023; Wan et al., 2023c; Peng et al., 2022; Mostafa, 2022; Ma et al., 2019). Mini-batch GNN training introduces additional variability in the gradient estimation, even compared to regular mini-batch DNN training. Specifically, in regular DNN training, the stochasticity in gradient estimation primarily arises from using mini-batches, which are random subsets of the training data, in each iteration. In mini-batch GNN training, the graph sampling used to form micro-batches is an additional source of stochasticity beyond the sampling of vertices in the mini-batch, because the training algorithm is not exposed to all the required dependent data in the graph.
@@
Motivation: What is the state of the art? Despite their different approaches and bottlenecks, full-graph and mini-batch training systems share the same goal: efficient and accurate training of Graph Neural Networks (GNNs). It is therefore important to establish which GNN training systems, be they full-graph or mini-batch systems, converge faster and to a higher accuracy under different scenarios. There is, however, no clear answer to this question in the literature. We performed an exhaustive review of the existing literature and found that prior works that propose systems in one class compare them with other systems in the same class, with limited or no performance evaluation of systems in the other class. We report a detailed discussion of the evidence in the literature in Section  2. Overall, the existing literature makes it difficult to get a clear picture of the state-of-the-art of the field.
Performance evaluation. In this paper, we address this open issue by performing a thorough performance comparison of representative state-of-the-art GNN training systems across the two classes. For full-graph training, we consider PipeGCN (Wan et al., 2022b), BNS-GCN (Wan et al., 2022a), and AdaQP (Wan et al., [n.d.]b). For mini-batch training, we consider DGL (Wang et al., 2019), DistDGL (Zheng et al., 2020), and Quiver (Tan et al., 2023), and three sampling algorithms: Neighborhood Sampling (Hamilton et al., 2017), ClusterGCN (Chiang et al., 2019), and GraphSaint (Zeng et al., 2019).
@@
Our results show that the mini-batch training systems are consistently faster than the full-graph training counterparts in reaching the same target accuracy across all datasets, GNN models, and hardware deployments we considered. These results are consistent with the evidence in the literature. Several papers that propose full-graph training systems use epoch time, which is the average time needed to execute one epoch, as the main metric to compare with mini-batch training systems, showing that the latter has a larger epoch time  (Wang et al., 2022; Wan et al., 2023b; Md et al., 2021; Chen et al., 2023). Our evaluation confirms these results. However, our results also highlight that time-to-accuracy is a better metric for comparison. Mini-batch training typically requires fewer epochs to converge because it updates the model parameters multiple times per epoch, once at each iteration, whereas full-graph training performs only one update per epoch. Therefore, even if mini-batch training systems perform more work per epoch than full-graph training ones to run multiple iterations, their overall time-to-accuracy is lower. Our empirical results show that this holds consistently across all systems, deployments, models, and datasets we consider. We report our performance results in Section 4 and support and generalize our empirical observations with an analysis in Section 6.
Performance vs. accuracy. Using a system with a longer time-to-accuracy can be justifiable if a model can converge to a higher accuracy. This is an important factor when navigating the performance trade-offs of using different GNN training systems.
@@
Our results show that, with proper hyperparameter tuning, mini-batch training can reach a similar accuracy as, if not higher accuracy than, full-graph training across different datasets and models. In our evaluation, we matched or exceeded all the test accuracy results for the GNN models we considered in the literature we reviewed. We observe that given a dataset and GNN model, hyperparameters that yield high accuracy for one training method do not work as well for the other. This may explain the contradicting results found in the literature and underline the importance of conducting separate hyperparameter tuning for the two training methods. These empirical findings also imply that filtering information during graph aggregation is not necessarily detrimental to accuracy, which is against the assumption made by works advocating full-graph training approaches. We discuss these results in Section 5.
@@
Lessons learned. Our results show that the mini-batch training systems we consider consistently achieve better performance than the full-graph training ones and have similar accuracy. We stress, however, that the goal of this study is not to establish the inherent superiority of one class of systems over the other. Our study is an empirical evaluation and the field is continuously evolving. Mini-batch training also has limitations, such as a slightly higher variance in accuracy between training runs and a high variance across different sampling algorithms.
Overall, our work highlights that existing algorithmic optimizations such as sampling or asynchrony can improve efficiency with no or minimal accuracy degradation, representing a promising research direction to achieve further performance and accuracy improvements. From a methodological perspective, our work highlights the importance of comparing GNN training systems across different classes and it also indicates a set of principles on how to perform this comparison.
@@ 2. Background and Motivation
This section provides the necessary background on GNN and GNN training systems to follow the rest of the paper. It then discusses the existing evidence in the literature.
@@ 2.1. Graph Neural Networks
Given a graph $G(V,E)$, where each node $v\in V$ is represented with a feature vector $h_{v}^{0}$ (input feature of vertex $v$) and the edges between any two vertices $u$ and $v$ is $e_{uv}\in E$. A GNN at $l^{th}$ layer performs the following computation to get the hidden features of a vector:
@@
where $h_{v}^{l}$ are the $l^{th}$ layer features for vertex $v$, $N$ represents the incoming neighbor vertices of vertex $v$, $W^{l}$ is the model weight matrix, $\Psi^{l}$ is any aggregation function and $\Phi^{l}$ is an update function.
@@
Proposes System/paper Main FG vs. MB Claims Experimental evidence Year Open Source Evaluation Setup
@@
ADGNN (Song et al., 2023a) FG higher ACC, FG better convergence 2023 Multi-host CPU, Multi-host single-GPU
G3 (Wan et al., 2023b) FG higher ACC, lower ET ACC, ET 2023 Multi-host multi-GPU
AdaQP (Wan et al., 2023c) MB information loss 2023 Multi-host multi-GPU
@@
DGCL (Cai et al., 2021) MB ACC loss 2021 Multi-host multi-GPU
@@
Full-graph and mini-batch training systems implement different data management pipelines for scalable multi-GPU training, which resulted in the emergence of two separate categories of systems. In this paper, we focus our evaluation on GPU-based training on multi-GPU systems, which is a common configuration for scalable GNN training. For a comprehensive and detailed discussion on distributed GNN training systems, see also (Shao et al., 2024; Lin et al., 2023).
Full-graph training systems. The key idea of full-graph training is to execute the GNN layers for all vertices in the graph, as shown in Eqn. 1. To scale to large graphs, distributed full-graph training systems partition the graph into multiple subgraphs such that each can fit into the memory of one GPU. They use a model-parallel approach to train across GPUs that requires exchanging vertex features across partitions. The main focus of work on full-graph training systems has been on reducing the cost of vertex feature communication. We now discuss existing full-graph GNN training systems work in terms of two design dimensions: training data management and algorithmic optimizations (see Figure 2).
Some full-graph training systems proposed optimizations that focused on the management of training data and do not introduce approximations that can potentially impact accuracy. If there are not enough GPUs to store each partition in a different GPU, it is necessary to transfer data between the host memory and GPU memory. Prior work such as NeuGraph (Ma et al., 2019), RoC (Jia et al., 2020), G3 (Wan et al., 2023b), and HongTu (Wang et al., 2023) have proposed techniques to optimize host-GPU communication. When there are enough GPUs to distribute all the partitions of the graph in GPU memory, training proceeds as depicted in Figure 1(a) (Wang et al., 2022; Wan et al., 2023b; Cai et al., 2021; Tripathy et al., 2020).
Some recent works improve performance by further introducing algorithmic optimizations that can impact accuracy, unlike the work we described previously. Sancus (Peng et al., 2022), PipeGCN (Wan et al., 2022b), and GNNPipe (Chen et al., 2023) introduce asynchrony and allow GPUs to operate on stale vertex features. This enables GPUs to overlap communication with computation instead of having to wait at each layer for fresh vertex features coming from other GPUs. An alternative research direction has been to use sampling. BNS-GCN samples boundary nodes with edges across partitions at each epoch and exchanges vertex features only for those vertices (Wan et al., 2022a). ADGNN proposes an aggregation-difference sampling algorithm (Song et al., 2023a). Finally, the AdaQP system proposed message quantization to reduce the communication cost, together with new partitioning algorithms (Wan et al., 2023c).
Figure 2. Classes of optimizations that impact performance (P) and accuracy (A), and representative systems evaluated in this work.
Mini-batch training systems. Mini-batch training systems break up the training dataset into mini-batches and train and update the model on one mini-batch at a time. They use data parallelism to scale to multiple GPUs. Each GPU computes the hidden features of a subset of the vertices in the mini-batch, which is called the micro-batch. In principle, to execute $k$ GNN layers without loss of information as shown in Eqn. 1, each GPU would need to load all the vertices in the $k$-hop neighborhood of the vertices in the micro-batch. The resulting subgraph, however, is often too large to be loaded into one GPU, inducing the so-called neighborhood explosion problem. Mini-batch training systems use sampling to load only a subset of the $k$-hop neighborhood. This sample-load-train pipeline is depicted in Figure 1(b). Compared to full-graph training, data parallelism eliminates the need to exchange hidden vertex features across partitions. The pipeline, however, introduces two different bottlenecks: sampling and data loading. We now discuss systems-level optimizations that mitigate these bottlenecks without impacting accuracy and then discuss the choice of sampling algorithms (see Fig 2).
Initial mini-batch systems ran CPU-based sampling, but this can represent a major performance bottleneck, motivating the need for GPU-based sampling (Jangda et al., 2021; Serafini and Guan, 2021). Several systems-level optimizations have been proposed to run sampling algorithms on GPUs efficiently (Jangda et al., 2021; Pandey et al., 2020; Wang et al., 2021; Gong et al., 2023; Cai et al., 2023).
In terms of management of the training data, loading the training data to the GPUs can be a significant performance bottleneck. To address it, prior work proposed caching training data in GPU memory, such as the input features (Lin et al., 2020) or the graph structure (Yang et al., 2022b; Zhang et al., 2023; Cai et al., 2023). In local GPU caching, each GPU only accesses its local cache (Lin et al., 2020; Yang et al., 2022b). In distributed GPU caching GPUs can directly access data cached in the main memory of other GPUs, which is faster than accessing data from the host memory when GPUs are connected through a fast bus like NVLink (Tan et al., 2023; Yang et al., 2022a; Song et al., 2023b; Cai et al., 2023). Mini-batch training systems can leverage GPU caching also when only a subset of the training dataset is cached (partial caching). They load data from the host memory in case of cache misses (Min et al., 2022; Yang et al., 2022b; Sun et al., 2023; Tan et al., 2023; Liu et al., 2023; Cai et al., 2023).
In terms of algorithmic optimizations that can impact accuracy, mini-batch training systems can run different sampling algorithms. The seminal work that proposed mini-batch GNN training, GraphSage, proposed a simple neighborhood sampling approach (Hamilton et al., 2017). Subsequent work proposed many different algorithms such as FastGCN (Chen et al., 2018a), ClusterGCN (Chiang et al., 2019), LADIES (Zou et al., 2019), or GraphSaint (Zeng et al., 2019).